I need help on this question. A culture of bacteria grew from 320 to 780 in 5 hours. If this growth continues at this rate then how long will it take the population to triple?
1
Expert's answer
2011-07-20T06:29:24-0400
Let`s assume that bacteria culture grow linearly.
x0 = 320, t = 5h, y1 = 780
y = αx0 t => α = y1/(x0 t) = 780/(320*5) = 0.4875
y = 3x0 = αx0 t => t = (3x0)/(αx0) = 3/α = 3/0.4875 = 80/13
So, the population will triple after 80/13 hours.